# DetailMaster: Can Your Text-to-Image Model Handle Long Prompts?

### v1

@@ Abstract
While recent text-to-image (T2I) models show impressive capabilities in synthesizing images from brief descriptions, their performance significantly degrades when confronted with long, detail-intensive prompts required in professional applications. We present DetailMaster, the first comprehensive benchmark specifically designed to evaluate T2I models’ systematical abilities to handle extended textual inputs that contain complex compositional requirements. Our benchmark introduces four critical evaluation dimensions: Character Attributes, Structured Character Locations, Multi-Dimensional Scene Attributes, and Explicit Spatial/Interactive Relationships. The benchmark comprises long and detail-rich prompts averaging 284.89 tokens, with high quality validated by expert annotators. Evaluation on 7 general-purpose and 5 long-prompt-optimized T2I models reveals critical performance limitations: state-of-the-art models achieve merely $\sim$50accuracy in key dimensions like attribute binding and spatial reasoning, while all models showing progressive performance degradation as prompt length increases. Our analysis highlights systemic failures in structural comprehension and detail overload handling, motivating future research into architectures with enhanced compositional reasoning. We open-source the dataset, data curation code, and evaluation tools to advance detail-rich T2I generation and enable broad applications that would otherwise be infeasible due to the lack of a dedicated benchmark.
11footnotetext: Equal contribution.
@@ 1 Introduction
Recently, text-to-image (T2I) generation has experienced remarkable development, achieving higher-quality image synthesis and expanding practical products dalle3 ; sd3 ; flux2024 ; sdxl . However, while text-conditioned control offers a convenient way to generate desired images, most current models are primarily optimized for short textual inputs. They often struggle with long and complex prompts (see Figure 1 and Appendix J for failure cases), which may consist of hundreds of tokens and contain rich details (e.g., character attributes, spatial relationships, and diverse scene properties).
@@
The limitations of text-to-image models in faithfully following long prompts can be attributed to four constraints: 1) Training data bias. The mainstream models are primarily trained on short prompts (e.g., COCO coco averaging 10.5 tokens, CC12M cc12m averaging 20.2 tokens), reinforcing a bias toward concise textual inputs. 2) Structural comprehension deficiency. Most text encoders show limitations in parsing hierarchical descriptions involving multiple objects, attributes, and their spatial relationships. They may incorrectly segment information and fall into attribute misalignment. 3) Detail overload. When prompts contain excessive descriptive details about a single subject, models tend to omit or distort specific characteristics. 4) Token length constraints. Most text encoders impose strict upper bounds on input tokens (e.g., CLIP’s 77-token limit clip ).
@@
The ability to faithfully adhere to long prompts is crucial in many practical scenarios where users have extensive requirements. Developing a robust benchmark for evaluating long-prompt-driven image generation is of significant research and practical value, across domains like interactive media, visual storytelling, scientific visualization, industrial prototyping, and so on aigc_survey ; storytelling ; aigc_design ; aigc_edu . Existing evaluations of text-to-image models have examined their capabilities across multiple dimensions, including image-text alignment, attribute binding, spatial relationships, and human preference. However, these evaluations predominantly rely on short prompts. A few existing benchmarks involve long-prompt scenarios DPG-Bench ; denseprompt , but they remain limited by constrained token lengths (typically around 100 tokens) and lack detailed attribute descriptions and scene context.
@@
To facilitate a more rigorous evaluation of text-to-image models in long-prompt scenarios, we introduce DetailMaster, a novel benchmark along with a robust evaluation protocol. Our benchmark is constructed based on four critical dimensions of prompt comprehension and adherence: 1) Character Attributes, 2) Structured Character Locations, 3) Multi-Dimensional Scene Attributes, and 4) Explicit Spatial or Interactive Relationships. The prompts in our benchmark are derived from existing detailed captioning data docci ; localized_narratives , which we further refine by augmenting missing visual details, resulting in an average token length of 284.89. To ensure high quality, human experts are engaged to evaluate our benchmark, confirming its high standard. Regarding the evaluation protocol, we develop specialized assessment mechanisms for each category of attributes, enabling systematic evaluation of models’ compositional T2I abilities in long-prompt scenarios.
Employing the DetailMaster benchmark, we investigate 7 general-purpose models, including state-of-the-art (SOTA) models FLUX flux2024 and GPT Image-1 gpt4o-image1 , along with 5 specialized models optimized for long prompts. The experimental results and analysis provide valuable insights into the critical limitations of current text-to-image techniques in handling detail-rich long prompts. For example, even the most advanced models achieve merely $\sim$50accuracy in generating correct character-level features in such scenario. Furthermore, the optimized models’ performance is heavily constrained by their backbone architectures. And there is a consistent degradation in prompt adherence as prompt length increases, highlighting fundamental challenges in handling long prompts.
@@
In summary, our contributions are three-fold: (1) We propose the DetailMaster benchmark, with a comprehensive evaluation dataset and a robust evaluation protocol to systematically assess T2I models’ prompt comprehension and adherence in long-prompt scenarios; (2) We conduct extensive comparisons on both popular text-to-image models and those specifically optimized for long prompts with DetailMaster, deriving rich insights for further improvements; (3) We open-source all data and code (https://github.com/modelscope/data-juicer/tree/DetailMaster) to facilitate future research in text-to-image generation for long, detail-rich prompts.
@@
Figure 1: Text-to-image errors in long prompt scenario (with FLUX.1-dev). Real source image (left), detailed caption/prompt (center), and generated image (right). Red text indicates failure points.
@@ 2 Related works
Text-to-image models.
Text-to-image generation aims to synthesize semantically aligned images conditioned on textual descriptions. Initial endeavors predominantly use Generative Adversarial Networks (GANs) gan , such as StackGAN stackgan and AttnGAN attngan , employing a generator-discriminator framework to produce images. Subsequent advancements introduce autoregressive models that treat image generation as a sequence prediction task. Models like DALL·E dalle and Parti parti employ transformer architectures to generate image tokens sequentially, conditioned on textual prompts. More recently, diffusion-based models diffusion have emerged as a powerful paradigm. These models, including DALL·E 2 dalle2 , Imagen imagen , and Stable Diffusion stablediffusion , operate by iteratively denoising random noise to produce coherent images. Beyond this, flow-based models such as Stable Diffusion 3.5 sd3 and FLUX flux2024 use flow matching flow_matching techniques to learn direct transformations from noise to images, reducing sampling complexity and improving efficiency. Furthermore, the proprietary models Gemini 2.0 Flash Gemini_image and GPT Image-1 gpt4o-image1 also show SOTA level performance.
@@
HRS-Comp hrs_bench 3004 620 1000 N/A N/A 2000 16.43
T2I-CompBench t2i_compbench 6000 2316 4000 N/A N/A 2000 12.65
DensePrompts denseprompt 7061 4913 N/A N/A N/A N/A 100.04
DPG-Bench DPG-Bench 1065 4286 5020 N/A 329 2593 83.91
@@
Early evaluations for text-to-image generation primarily use datasets such as CUB Birds cub_bird and Oxford Flowers oxford_flower , which have limited diversity and simple prompts. Recently, benchmarks like DrawBench imagen , PartiPrompts parti , HRS-Bench hrs_bench and T2I-CompBench t2i_compbench introduce diverse sets of prompts aimed at evaluating compositional generation abilities, including object presence, attribute binding, and spatial relationships. To precisely measure the alignment between generated images and every detail within the prompts, benchmarks such as TIFA tifa and Gecko gecko decompose prompts into elemental components and generate corresponding questions to assess model fidelity. Moreover, benchmarks like GenAI-Bench genai_bench , RichHF18K richhf and HPS v2 hpsv2 involve human annotators to rank image-text pairs, aiming to train preference-aligned evaluators that reflect human judgment. However, these benchmarks predominantly focus on short prompts. They overlook the long and complex prompts, which can significantly affect the generation with intricate elements and details, such as increasing the risk of attribute misalignment.
Text-to-image models for long prompt.
@@
To enhance text-to-image models’ ability to follow long prompts, some studies employ large language models (LLMs) as text encoders to improve semantic extraction. Representative models such as Imagen imagen , DeepFloyd IF DeepFloydIF , and advanced approaches like Stable Diffusion 3.5 sd3 and FLUX flux2024 adopt frozen T5 t5 as their text encoder, strengthening text-image alignment. Beyond this, ParaDiffusion paradiffusion uses a frozen Llama V2 llama2 as its text encoder and constructs a training dataset for long prompts to enhance comprehension. LLM4GEN denseprompt introduces an adapter to fuse features from multiple text encoders (e.g., T5 and CLIP) and proposes a specialized loss to penalize attribute mismatch. ELLA DPG-Bench incorporates a trainable mapper after the LLM-based text encoder to improve information extraction from long prompts. LongAlign longalign decomposes long prompts into individual sentences for separate encoding, while also enabling preference optimization for long-text alignment. LLM Blueprint Llm_blueprint extracts object details from long prompts using an LLM, then employs layout-to-image generation and an iterative refinement scheme for image synthesis. However, due to the scarcity of benchmarks for long-prompt-driven image generation, these methods lack fair comparison. Thus, we propose the DetailMaster benchmark to systematically assess their performance.
Benchmarks for long-prompt-driven text-to-image generation.
@@
Current benchmarks for long-prompt-driven image generation remain scarce. DensePrompts denseprompt collects 100 detailed web images, using GPT-4V gpt4v to generate attribute-rich descriptions and expand the length. DPG-Bench DPG-Bench aggregates multi-short-prompt annotations from existing datasets and employs GPT-4 gpt4 to synthesize them into long descriptions. However, these benchmarks exhibit critical limitations: 1) oversimplified evaluation, relying on CLIP Score clipscore for text-image alignment or binary feature verification via multimodal large language model (MLLMs), which fails to assess fine-grained and compositional accuracy; 2) constrained prompt length and details, deriving prompts through coarse LLM extraction and failing to match real-world long-prompt complexity. To address these issues, we propose DetailMaster, a novel benchmark featuring longer prompts along with a fine-grained evaluation protocol to rigorously assess text-to-image generation under complex long-prompt scenarios. As evidenced in Table 1, our benchmark advances prior works through extended prompts, enlarged attribute sets, and more comprehensive evaluation objectives.
@@ 3.1.1 Overview
To better simulate the challenges in long-prompt scenarios, our dataset requires prompts of sufficient length and rich details. Hence, we leverage existing human-annotated detailed captioning datasets, where each sample comprises an image paired with fine-grained textual descriptions, including DOCCI (15K samples) docci and Localized Narratives (Flickr30k flickr30k and COCO subsets, 149K samples) localized_narratives , and expand their captions. While these datasets already provide precise object attribute descriptions with an average token length of 68.4, their coverage is still incomplete. To enhance image description quality and obtain optimal prompts, we design a robust and fine-grained attribute extraction pipeline that systematically captures four categories of key features: 1) Character Attributes, where we categorize main characters (objects, animals, persons, and others) and analyze them via class-specific attribute extraction protocols; 2) Structured Character Locations, where we employ a nine grid-based spatial partitioning scheme to assign precise positions; 3) Multi-Dimensional Scene Attributes, decomposing scenes hierarchically into background, lighting conditions, and stylistic elements; and 4) Explicit Spatial/Interactive Relationships, annotating character pairs with geometric (e.g., “dog behind chair”) and dynamic interactions (e.g., “holding”). These extracted attributes are then integrated into the original captions using LLM-based expansion, yielding richly detailed prompts with an average token length of 284.89. Beyond prompt enhancement, they also form the foundation for constructing specialized evaluation metrics. The overview of the data generation process is shown in Figure 2. To enhance reusability, our data construction process is also implemented as data processing operators and configurable data recipes within Data-Juicer dj ; djsandbox .
@@ 3.1.2 Attribute extraction pipeline
Main character identification.
We first need to know what characters are present in each sample. Specifically, we provide an MLLM with both the source image and the original caption to detect main characters within the image. We then apply sample filtering to retain only instances containing more than 4 main characters, ensuring as many high-quality features as possible in the final prompt to better evaluate models’ structural comprehension and their ability to handle detail overload.
Character localization.
For character localization, we detecting each main character using both an MLLM and the open-set object detector YOLOE-11L yoloe to generate two bounding box predictions. We then establish a validation where the two boxes with an Intersection over Union (IoU) score at least 0.7 are considered mutually confirmatory, with the larger one selected as the valid bounding box. For cases with IoU less than 0.7, we crop sub-images of each box and employ BLIP blip to compute image-text matching scores between each sub-image and the character name. If both scores fall below 0.4, we discard both detections; otherwise, we retain the higher-scoring box. Finally, we transform validated bounding boxes into structured spatial descriptions using the nine-region partitioning scheme detailed in Appendix A. They are recorded as “Structured Character Locations”.
Character attribute extraction.
For character attribute extraction, we implement a hierarchical pipeline. Initially, each main character is classified by an LLM into one of four predefined categories: object, animal, person, or other. Subsequently, category-specific instructions are applied to guide the LLM in extracting relevant character attributes from the original captions, such as material and shape for objects, color and movement patterns for animals, and clothing and age for persons, thereby generating preliminary attribute lists. To enhance attribute completeness and simulate detail overload, we then implement a multimodal refinement stage. For characters with valid bounding box information, we extract corresponding sub-images and ask an MLLM to supplement the text-derived attribute lists with visually-grounded features. The synthesized output from both textual and visual analysis constitutes the final “Character Attributes” for each character.
Scene attribute and entity relationship detection.
For scene attributes extraction, we employ an MLLM to jointly analyze both the source image and the original caption, characterizing background composition, lighting conditions, and stylistic elements of each sample. These features are recorded as “Scene Attributes”. Regarding entity relationship extraction, we input each sample’s source image, original caption, and all detected characters, into the MLLM. The model then performs relationship parsing to identify and extract all discernible spatial configurations and interactive relationships between entities, which are finally recorded as “Explicit Spatial/Interactive Relationships.”

### v2

@@ Abstract
While recent text-to-image (T2I) models show impressive capabilities in synthesizing images from brief descriptions, their performance significantly degrades when confronted with long, detail-intensive prompts required in professional applications. We present DetailMaster, the first comprehensive benchmark specifically designed to evaluate T2I models’ systematic abilities to handle extended textual inputs that contain complex compositional requirements. Our benchmark introduces four critical evaluation dimensions: Character Attributes, Structured Character Locations, Multi-Dimensional Scene Attributes, and Spatial/Interactive Relationships. The benchmark comprises long and detail-rich prompts averaging 284.89 tokens, with high quality validated by expert annotators. Evaluation on 7 general-purpose and 5 long-prompt-optimized T2I models reveals critical performance limitations: state-of-the-art models achieve merely $\sim$50% accuracy in key dimensions like attribute binding and spatial reasoning, while all models showing progressive performance degradation as prompt length increases. Our analysis reveals fundamental limitations in compositional reasoning, demonstrating that current encoders flatten complex grammatical structures and that diffusion models suffer from attribute leakage under detail-intensive conditions. We open-source our dataset, data curation code, and evaluation tools to advance detail-rich T2I generation and enable applications previously hindered by the lack of a dedicated benchmark.
@@ 1 Introduction
While recent advancements in text-to-image (T2I) generation have produced high-fidelity images and practical applications (Betker et al., 2023; Esser et al., 2024; Labs, 2024; Podell et al., 2023), current models remain primarily optimized for short textual inputs. They often struggle with long, complex prompts (see Figure 1 and Appendix P for failure cases), which may consist of hundreds of tokens and contain rich details (e.g., character attributes, spatial relationships, and scene properties).
@@
The limitations of T2I models in faithfully following long prompts can be attributed to four constraints: 1) Training data bias. Models are primarily trained on short prompts (e.g., COCO (Lin et al., 2014) averaging 10.5 tokens, CC12M (Changpinyo et al., 2021) averaging 20.2 tokens), which reinforces a preference for concise inputs. 2) Structural comprehension deficiency. Most text encoders fail to adequately parse hierarchical descriptions involving multiple objects, attributes, and their spatial relationships, which results in incorrect information segmentation and attribute misalignment. 3) Detail overload. When prompts contain excessive descriptive details on a single subject, models tend to omit or distort key details. 4) Token length constraints. Most encoders impose strict upper bounds on input tokens (e.g., CLIP’s 77-token limit (Radford et al., 2021)).
@@
Faithfully interpreting long prompts is a critical capability in many practical scenarios where users have extensive requirements, such as interactive media, visual storytelling, scientific visualization, industrial prototyping, and so on (Cao et al., 2025; Huang et al., 2016; Wu et al., 2024; Xing et al., 2023). Existing evaluations of T2I models have examined their capabilities across multiple dimensions, including image-text alignment, attribute binding, and human preference. However, these evaluations predominantly rely on short prompts. Only a few existing benchmarks involve long-prompt scenarios (Hu et al., 2024; Liu et al., 2025), but they remain limited by constrained token lengths (typically around 100 tokens) and lack detailed attribute descriptions and scene context.
@@
To rigorously evaluate T2I models on long prompts, we introduce DetailMaster, a novel benchmark along with a robust evaluation protocol. This benchmark assesses prompt adherence across four critical dimensions, specifically designed to target two challenges: structural comprehension deficiency (via Character Locations and Spatial/Interactive Relationships) and detail overload (via Character Attributes and Scene Attributes). The prompts in our benchmark are derived from existing detailed captions (Onoe et al., 2024; Pont-Tuset et al., 2020), which we further refine by augmenting missing details, resulting in an average token length of 284.89. To ensure high quality, human experts are engaged to evaluate our benchmark, confirming its high standard. For the evaluation protocol, we develop specialized assessment mechanisms for each category of attributes, enabling systematic evaluation of models’ compositional T2I abilities in long-prompt scenarios.
Using our DetailMaster benchmark, we investigate 7 general-purpose models, including state-of-the-art (SOTA) models FLUX (Labs, 2024) and GPT Image-1 (OpenAI, 2025), along with 5 specialized models optimized for long prompts. While our analysis validates enhanced competencies in recent models over older baselines, it also uncovers critical failure points. We identify a consistent struggle with fine-grained compositional elements, including character attributes, spatial positioning, and entity relationships. A statistical analysis confirms a clear difficulty hierarchy, with Character Locations and Person Attributes proving the most challenging. Furthermore, we find that performance gains are not merely a function of longer context windows; training on detail-rich, longer prompts is essential. Furthermore, the optimized models’ performance is heavily constrained by their backbone architectures. And there is a consistent degradation in prompt adherence as prompt length increases, highlighting fundamental challenges in handling long prompts.
@@
In summary, our contributions are three-fold: 1) We propose the DetailMaster benchmark, featuring a comprehensive dataset and a robust evaluation protocol to systematically assess the prompt adherence of T2I models in long-prompt scenarios. 2) We conduct extensive comparisons on both popular T2I models and those specifically optimized for long prompts, yielding several insights: current models fail at complex compositional tasks and show a degradation in adherence as prompt length increases; progress in long-prompt adherence is driven less by expanded context windows and more by crucial factors such as long-prompt training and iterative decomposition. 3) We open-source all data and code (https://github.com/modelscope/data-juicer/tree/DetailMaster) to facilitate future research in text-to-image generation for long, detail-rich prompts.
@@
Figure 1: Text-to-image errors in long prompt scenario (with FLUX.1-dev). Real source image (left), detailed caption/prompt (middle), and generated image (right). Red text indicates failure points.
@@ 2 Related Works
Text-to-image models. Text-to-image (T2I) generation aims to synthesize semantically aligned images conditioned on textual descriptions. Initial endeavors use Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), employing a generator-discriminator framework to produce images. Subsequent advancements introduce autoregressive models that treat image generation as a sequence prediction task, including DALL·E (Ramesh et al., 2021) and Parti (Yu et al., 2022). More recently, diffusion-based models (Ho et al., 2020) have emerged as a powerful paradigm. These models, including DALL·E 2 (Ramesh et al., 2022), Imagen (Saharia et al., 2022), and Stable Diffusion (Rombach et al., 2022), operate by iteratively denoising to produce coherent images. Beyond this, flow-based models such as Stable Diffusion 3.5 (Esser et al., 2024) and FLUX (Labs, 2024) use flow matching (Lipman et al., 2022) techniques to learn direct transformations from noise to images, reducing sampling complexity and improving efficiency. Furthermore, the proprietary models Gemini 2.0 Flash (Google, 2025) and GPT Image-1 (OpenAI, 2025) also show SOTA level performance.
@@
HRS-Comp (Bakr et al., 2023) 3004 620 1000 N/A N/A 2000 16.43
T2I-CompBench (Huang et al., 2023) 6000 2316 4000 N/A N/A 2000 12.65
DensePrompts (Liu et al., 2025) 7061 4913 N/A N/A N/A N/A 100.04
DPG-Bench (Hu et al., 2024) 1065 4286 5020 N/A 329 2593 83.91
@@
Benchmarks for text-to-image generation. Early evaluations for T2I generation primarily use datasets such as CUB Birds (Wah et al., 2011) and Oxford Flowers (Nilsback & Zisserman, 2008), which have limited diversity and simple prompts. Recently, benchmarks such as DrawBench (Saharia et al., 2022), HRS-Bench (Bakr et al., 2023) and T2I-CompBench (Huang et al., 2023) introduce prompts aimed at compositional generation, including object presence, attribute binding, and spatial relationships. To measure the alignment between generated images and every detail within the prompts, benchmarks such as TIFA (Hu et al., 2023) and Gecko (Wiles et al., 2024) decompose prompts into elemental components and generate corresponding questions to assess model fidelity. Moreover, benchmarks such as GenAI-Bench (Li et al., 2024a), RichHF18K (Liang et al., 2024) and HPS v2 (Wu et al., 2023c) aim to train preference-aligned evaluators that reflect human judgment. However, these benchmarks predominantly focus on short prompts, neglecting the challenges of longer, complex inputs that are prone to issues like attribute misalignment.
@@
T2I models for long prompt. To enhance T2I models’ ability to follow long prompts (those exceeding the 77-token limit of the CLIP text encoder), some studies employ large language models (LLMs) as text encoders to improve semantic extraction. Representative models such as Imagen (Saharia et al., 2022), DeepFloyd IF (at StabilityAI, 2023), and advanced approaches like Stable Diffusion 3.5 (Esser et al., 2024) and FLUX (Labs, 2024) adopt frozen T5 (Raffel et al., 2020) as their text encoder, strengthening text-image alignment. Beyond this, ParaDiffusion (Wu et al., 2023a) uses a frozen Llama V2 (Touvron et al., 2023) as its text encoder and constructs a training dataset of long prompts to enhance comprehension. LLM4GEN (Liu et al., 2025) introduces an adapter to fuse features from multiple text encoders (e.g., T5 and CLIP) and proposes a specialized loss to penalize attribute mismatch. ELLA (Hu et al., 2024) incorporates a trainable mapper after the LLM-based text encoder to improve information extraction from long prompts. LongAlign (Liu et al., 2024) decomposes long prompts into individual sentences for separate encoding, while also enabling preference optimization for long-text alignment. LLM Blueprint (Gani et al., 2023) extracts object details from long prompts using an LLM, then employs layout-to-image generation and an iterative refinement scheme for image synthesis. However, due to the scarcity of benchmarks for long-prompt-driven image generation, these methods lack fair comparison.
@@
Benchmarks for long-prompt-driven T2I generation. Current benchmarks for long-prompt-driven image generation remain scarce. DensePrompts (Liu et al., 2025) collects 100 detailed web images, using GPT-4V (Achiam et al., 2023b) to generate attribute-rich descriptions. DPG-Bench (Hu et al., 2024) aggregates multi-short-prompt annotations from existing datasets and employs GPT-4 (Achiam et al., 2023a) to synthesize them into long descriptions. However, these benchmarks exhibit critical limitations: 1) oversimplified evaluation, relying on CLIP Score (Hessel et al., 2021) or binary feature verification via multimodal large language model (MLLMs), which fail to assess fine-grained and compositional accuracy; 2) constrained prompt length and details, deriving prompts through LLM extraction and failing to match real-world long-prompt complexity. To address these issues, we propose DetailMaster, a novel benchmark featuring longer prompts along with a fine-grained evaluation protocol to assess T2I generation under complex long-prompt scenarios. As evidenced in Table 1, our benchmark advances prior works through extended prompts, enlarged attribute sets, and more comprehensive evaluation objectives.
@@ 3.1.1 Overview
To better simulate the challenges in long-prompt scenarios, we require prompts of sufficient length and rich details. Hence, we leverage existing human-annotated detailed captioning datasets, where each sample comprises an image paired with fine-grained textual descriptions, including DOCCI (Onoe et al., 2024) and Localized Narratives (Flickr30k (Young et al., 2014) and COCO subsets) (Pont-Tuset et al., 2020). While these datasets already provide precise object descriptions with an average token length of 68.4, their coverage is still incomplete. To enhance the image description quality and obtain optimal prompts, we design a robust and fine-grained attribute extraction pipeline that systematically captures four categories of key features: 1) Character Attributes, where we categorize main characters (objects, animals, persons, and others) and analyze them via class-specific attribute extraction protocols; 2) Structured Character Locations, where we employ a nine grid-based spatial partitioning scheme to assign precise positions; 3) Multi-Dimensional Scene Attributes, decomposing scenes into background, lighting conditions, and stylistic elements; and 4) Spatial/Interactive Relationships, annotating character pairs with geometric (e.g., “dog behind chair”) and dynamic interactions (e.g., “holding”). These extracted attributes are then integrated into the original captions using LLM-based expansion, yielding richly detailed prompts with an average token length of 284.89. The overview of the data generation process is shown in Figure 2. To enhance reusability, our data construction process is also implemented as data processing operators and configurable data recipes within Data-Juicer (Chen et al., 2025a; b).
@@ 3.1.2 Attribute Extraction Pipeline
Main character identification. We first need to know what characters are present in each sample. Specifically, we provide an MLLM with both the image and the caption to detect main characters. We then apply sample filtering to retain only instances containing more than four main characters, ensuring as many high-quality features as possible in the final prompt (see Figure 2 ①).
Character localization. For character localization, we generate two bounding box proposals using an MLLM and the open-set detector YOLOE-11L (Wang et al., 2025). We then validate these proposals through a two-stage process. First, if the boxes have an Intersection over Union (IoU) of at least 0.7, we select the larger one. Otherwise, we employ BLIP (Li et al., 2022b) to score each cropped box against the character’s name; we then retain the higher-scoring box unless both scores fall below 0.4 (see Figure 2 ②). Finally, the validated bounding boxes are converted into structured spatial descriptions using the nine-region partitioning scheme detailed in Appendix B.
Character attribute extraction. For character attributes, we implement a hierarchical pipeline. First, an LLM performs text-based extraction by classifying characters (object, animal, person, other) and using category-specific prompts to derive attributes from captions (e.g., material for objects, clothing for persons). Subsequently, a multimodal refinement stage enhances attribute completeness and simulate detail overload. For characters with valid bounding box information, we crop corresponding sub-images and ask an MLLM to supplement the text-derived attribute lists with visually-grounded features (see Figure 2 ③).
Scene attribute extraction. For scene attributes, we employ an MLLM to jointly analyze the source image and the caption for each sample, identifying its background composition, lighting conditions, and stylistic elements (see Figure 2 ④).
Entity relationship detection. For entity relationships, we input the source image, caption, and all detected characters of each sample into the MLLM. The model then performs relationship parsing to extract all discernible spatial and interactive relationships between entities (see Figure 2 ⑤).